OpenStudy (explainitlikeimfive):

Find the domain, period, range, and amplitude of the cosine function. y = -6cos4x

OpenStudy (explainitlikeimfive):

I have 4 choices, but I think it's this one.

OpenStudy (explainitlikeimfive):

domain = all real numbers; period = 1/2 ; range:-6<y<6 ; amplitude = -6

OpenStudy (anonymous):

Domain will always be \[(-\infty, \infty)\]

ganeshie8 (ganeshie8):

amplitude can never be negative

ganeshie8 (ganeshie8):

and you missed a pi in the period :)

OpenStudy (explainitlikeimfive):

omain = all real numbers; period = 1/2 ; range:-6<y<6 ; amplitude = 6 This is another choice. And none of them have pie.

OpenStudy (anonymous):

\[Period \frac{ 2\pi }{ (bx) }\]

OpenStudy (theeric):

I agree with the others. Here's a good site for identifying the parts! http://www.regentsprep.org/Regents/math/algtrig/ATT7/sinusoidal.htm Is says the amplitude will be the absolute value of what's multiplying the \(cos()\) function. Sorry to mislead you earlier!

OpenStudy (anonymous):

Yes what @theEric said about the abs calue on it and also amp is the number next to the sin so is it -6 or 6, there's a differnce in those two answers don't get tripped up

ganeshie8 (ganeshie8):

:) you must have the pi for the period cos(4x) runs 4 times faster than cos(x) since cos(x) period is 2pi, cos(4x) period wud be 2pi/4 = pi/2

OpenStudy (theeric):

The period should be \(\LARGE \frac{2\pi}{4}=\frac{\pi}{2}\)...

OpenStudy (theeric):

I'm guessing its something weird with the problem and its choices...

OpenStudy (anonymous):

How do you get your equation to be so big @theEric

OpenStudy (theeric):

I had a Pennsylvania standardized test not have a \(\pm\) option when solving for \(x\) under a radical...

OpenStudy (explainitlikeimfive):

Heres a screenshot of the exact problem.

ganeshie8 (ganeshie8):

go with B @ExplainItLikeImFive

OpenStudy (anonymous):

Well @ExplainItLikeImFive If you know teh amplitude is -6 and the domain is ARN and there's only one option with those two specifications, there's no need to calculate the rest of them if you already know how to find them (test taking tip)

OpenStudy (theeric):

When you're in the `\[ \]` or `\( \)`, use `\large`, `\Large`, `\LARGE`, `\huge`, or `\Huge`. They get bigger as you go along.

OpenStudy (anonymous):

Go with C the amplitude is -6 not 6 @ganeshie8

OpenStudy (anonymous):

Thanks!

ganeshie8 (ganeshie8):

amplitude is the distance from average value so u have to take the absolute value of the quantity next to cos amplitude = |-6| = 6

OpenStudy (theeric):

Amplitude is \(\left|-6\right|=6\), conventionally.

OpenStudy (anonymous):

My bad, ha, abs values always get me x.x

OpenStudy (theeric):

@doulikepiecauseidont and you can use `\tiny`, `\Tiny`, `\small`, and `\normalsize` too.

OpenStudy (theeric):

Haha!

OpenStudy (explainitlikeimfive):

Thanks to you all.

ganeshie8 (ganeshie8):

np :)

OpenStudy (explainitlikeimfive):

My only regret is I cant give all of you a medal.

OpenStudy (anonymous):

\[\Huge wow this is cool \]

OpenStudy (anonymous):

How about we al;l give each other one?

OpenStudy (theeric):

I know the feeling. They got to it first, so that helps narrow it down!

OpenStudy (theeric):

Alright, we can try that! :)

OpenStudy (explainitlikeimfive):

I christen @ganeshie8

OpenStudy (theeric):

Alright! And he got me and I got doulikepiecauseidont.

OpenStudy (theeric):

ganeshie must get ExplainItLikeImFive.

ganeshie8 (ganeshie8):

thank you :) @doulikepiecauseidont give it to @ExplainItLikeImFive then we all are even :)

OpenStudy (anonymous):

I did give it to him

OpenStudy (anonymous):

Someone already got you, thats why you hahve two

OpenStudy (theeric):

I clicked the wrong one the first time, but then I got ExplainItLikeImFive , now it's still uneven...

OpenStudy (anonymous):

Well i picked @ganeshie8 and so did @ExplainItLikeImFive

OpenStudy (explainitlikeimfive):

So I switch mine to @doulikepiecauseidont

OpenStudy (anonymous):

Thats why it was two but ok lets get this

OpenStudy (anonymous):

yea thats it

OpenStudy (explainitlikeimfive):

Alright, playing field leveled.

OpenStudy (theeric):

1's across the board. Success!